# Nearly Minimax Optimal Regret for Multinomial Logistic Bandit

### v1

@@ C.3 Proofs of Lemmas for Theorem 1
$=\frac{v_{0}K\left(\exp\left(x^{\top}\mathbf{w}_{V}\right)-\exp\left(\hat{x}^{\top}\mathbf{w}_{V}\right)\right)}{\left(v_{0}+K\exp\left(x^{\top}\mathbf{w}_{V}\right)\right)\left(v_{0}+\exp\left(\hat{x}^{\top}\mathbf{w}_{V}\right)\right)}$
$\geq\frac{v_{0}K\left(\exp\left(x^{\top}\mathbf{w}_{V}\right)-\exp\left(\hat{x}^{\top}\mathbf{w}_{V}\right)\right)}{(v_{0}+Ke)^{2}},$ (C.5)
where the inequality holds since $\max\left\{\exp\left(x^{\top}\mathbf{w}_{V}\right),\exp\left(\hat{x}^{\top}\mathbf{w}_{V}\right)\right\}\leq e$. To further bound the right-hand side of  (C.5), we use the fact that $21+a\leq e^{a}\leq 1+a+a^{2}/$ 2 for all $a\in[0,1]$, which can be easily shown by Taylor expansion. Thus, we get
@@
where the last inequality holds because $d\epsilon^{2}\leq\delta\epsilon/\sqrt{d}$ when $\epsilon\in(0,1/d\sqrt{d})$. This concludes the proof. ∎
C.3.2 Proof of Lemma C.2
Proof of Lemma C.2.
Fix a round $t$, an assortment $\tilde{S}_{t}$, and $\tilde{U}_{t}$. Let $U=\tilde{U}_{t}$. Define $m_{j}(\tilde{S}_{t}):=\sum_{x_{U}\in\tilde{S}_{t}}\mathbf{1}\{j\in U\}/K$. Let $\{p_{i}\}_{i\in\tilde{S}_{t}\cup\{0\}}$ and $\{q_{i}\}_{i\in\tilde{S}_{t}\cup\{0\}}$ be the probabilities of choosing item $i$ under parameterization $\mathbf{w}_{V}$ and $\mathbf{w}_{V\cup\{j\}}$, respectively. Then, we have
$\operatorname{KL}\left(P_{V}(\cdot|\tilde{S}_{t})\|P_{V\cup\{j\}}(\cdot|\tilde{S}_{t})\right)=\sum_{i\in\tilde{S}_{t}\cup\{0\}}p_{i}\log\frac{p_{i}}{q_{i}}\leq\sum_{i\in\tilde{S}_{t}\cup\{0\}}p_{i}\frac{p_{i}-q_{i}}{q_{i}}\leq\sum_{i\in\tilde{S}_{t}\cup\{0\}}\frac{(p_{i}-q_{i})^{2}}{q_{i}},$
@@
$\leq C_{0}\cdot\frac{v_{0}K^{2}}{(v_{0}+K)^{3}}\cdot\frac{m_{j}(\tilde{S}_{t})\epsilon^{2}}{d},$ (C.6)
@@
$\sum_{i\in\tilde{S}_{t},j\notin U}\frac{(p_{i}-q_{i})^{2}}{q_{i}}=0,$ (C.7)
@@
Recall that for any $i\in\tilde{S}_{t}$, $q_{i}\geq\frac{e^{-1}}{v_{0}+Ke}$ Then, for any $i\in\tilde{S}_{t}$ corresponding to $x_{i}=\hat{x}$ and $j\in U$, we have
@@
$\leq C_{1}\cdot\frac{v_{0}^{2}K}{(v_{0}+K)^{3}}\cdot\frac{m_{j}(\tilde{S}_{t})\epsilon^{2}}{d},$ (C.8)
@@
Combining (C.6), (C.7), and (C.8), we derive that
@@
$\operatorname{KL}\left(P_{V}\|Q_{V\cup\{j\}}\right)$ $=\sum_{t=1}^{T}\mathbb{E}_{V}\left[\operatorname{KL}\left(P_{V}(\cdot|\tilde{S}_{t})\|P_{V\cup\{j\}}(\cdot|\tilde{S}_{t})\right)\right]$
@@
where the first equality holds by the chain rule of relative entropy (cf. Exercise 14.11 of Lattimore and Szepesvári [17]). This concludes the proof. ∎
Appendix D Proof of Theorem 2
@@
D.1 Main Proof of Theorem 2
Before presenting the proof, we introduce useful lemmas, whose proof can be found in Appendix D.2. Lemma D.1 shows the optimistic utility for the centralized context vectors
Lemma D.1.
Let $\alpha_{ti}=x_{ti}^{\top}\mathbf{w}_{t}+\beta_{t}\|x_{ti}\|_{H_{t}^{-1}}$. If $\mathbf{w}^{\star}\in\mathcal{C}_{t}(\delta)$, then we have
@@ D.1 Main Proof of Theorem 2
Lemma D.3 is an elliptical potential lemma that improves upon the one presented in Lemma 10 of Goyal and Perivier [15] in terms of $K$.
Lemma D.2.
@@
Moreover, we provide a tighter bound for the second derivative of the revenue than that presented in Lemma 12 of Goyal and Perivier [15].
Lemma D.3.
Define $Q:\mathbb{R}^{K}\rightarrow\mathbb{R}$, such that for any $\mathbf{u}=(u_{1},\dots,u_{K})\in\mathbb{R}^{K}$, $Q(\mathbf{u})=\sum_{i=1}^{K}\frac{\exp(u_{i})}{\bar{v}_{0}+\sum_{k=1}^{K}\exp(u_{k})}$. Let $p_{i}(\mathbf{u})=\frac{\exp(u_{i})}{\bar{v}_{0}+\sum_{k=1}^{K}\exp(u_{k})}$. Then, for all $i\in[K]$, we have
@@
First, we express the regret as follows:
$\textup{{Reg}}_{T}(\mathbf{w}^{\star})$ $=\sum_{t=1}^{T}R_{t}(S_{t}^{\star},\mathbf{w}^{\star})-R_{t}(S_{t},\mathbf{w}^{\star})=\sum_{t=1}^{T}\left[\sum_{i\in S_{t}^{\star}}p_{t}(i|S_{t}^{\star},\mathbf{w}^{\star})-\sum_{i\in S_{t}}p_{t}(i|S_{t},\mathbf{w}^{\star})\right]$
@@
$\leq\sum_{t=1}^{T}\left[\frac{\sum_{i\in S_{t}}\exp(\alpha_{ti})}{v_{0}+\sum_{j\in S_{t}}\exp(\alpha_{tj})}-\frac{\sum_{i\in S_{t}}\exp(x_{ti}^{\top}\mathbf{w}^{\star})}{v_{0}+\sum_{j\in S_{t}}\exp(x_{tj}^{\top}\mathbf{w}^{\star})}\right]$
$=\sum_{t=1}^{T}\tilde{R}_{t}(S_{t})-R_{t}(S_{t},\mathbf{w}^{\star}),$
where the first inequality holds by Lemma D.1, and the last inequality holds by the assortment selection of Algorithm 1.
Now, we define $Q:\mathbb{R}^{K}\rightarrow\mathbb{R}$, such that for all $\mathbf{u}=(u_{1},\dots,u_{K})\in\mathbb{R}^{K}$, $Q(\mathbf{u})=\sum_{i=1}^{K}\frac{\exp(u_{i})}{v_{0}+\sum_{j=1}^{K}\exp(u_{j})}$. Noting that $S_{t}$ always contains $K$ elements since the expected revenue is an increasing function in the uniform reward setting, we write $S_{t}=\{i_{1},\dots,i_{K}\}$. Moreover, for all $t\geq 1$, let $\mathbf{u}_{t}=(u_{ti_{1}},\dots u_{ti_{K}})^{\top}=(\alpha_{ti_{1}},\dots,\alpha_{ti_{K}})^{\top}$ and $\mathbf{u}^{\star}_{t}=(u_{ti_{1}}^{\star},\dots u_{ti_{K}}^{\star})^{\top}=(x_{ti_{1}}^{\top}\mathbf{w}^{\star},\dots,x_{ti_{K}}^{\top}\mathbf{w}^{\star})^{\top}$. Then, by a second order Taylor expansion, we have
@@
$=\underbrace{\sum_{t=1}^{T}\nabla Q(\mathbf{u}_{t}^{\star})^{\top}(\mathbf{u}_{t}-\mathbf{u}^{\star}_{t})}_{\texttt{(A)}}+\underbrace{\frac{1}{2}\sum_{t=1}^{T}(\mathbf{u}_{t}-\mathbf{u}^{\star}_{t})^{\top}\nabla^{2}Q(\bar{\mathbf{u}}_{t})(\mathbf{u}_{t}-\mathbf{u}^{\star}_{t})}_{\texttt{(B)}},$
@@
$\leq 2\beta_{T}(\delta)\sum_{t=1}^{T}\sum_{i\in S_{t}}p_{t}(i|S_{t},\mathbf{w}^{\star})p_{t}(0|S_{t},\mathbf{w}^{\star})\|x_{ti}\|_{H_{t}^{-1}},$
where the first inequality holds by Lemma D.1, and the last inequality holds because $\beta_{t}$ is increasing for $t\in[T]$.
@@
$\leq\sum_{t=1}^{T}\sum_{i\in S_{t}}\sum_{j\in S_{t},j\neq i}|u_{ti}-u_{ti}^{\star}|p_{i}(\bar{\mathbf{u}}_{t})p_{j}(\bar{\mathbf{u}}_{t})|u_{tj}-u_{tj}^{\star}|+\frac{3}{2}\sum_{t=1}^{T}\sum_{i\in S_{t}}(u_{ti}-u_{ti}^{\star})^{2}p_{i}(\bar{\mathbf{u}}_{t}),$ (D.1)
where the inequality is by Lemma D.3. To bound the first term in (D.1), by applying the AM-GM inequality, we get
@@
$\leq\sum_{t=1}^{T}\sum_{i\in S_{t}}(u_{ti}-u_{ti}^{\star})^{2}p_{i}(\bar{\mathbf{u}}_{t}).$ (D.2)
By plugging (D.2) into (D.1), we have
@@
$\leq 10\beta_{T}(\delta)^{2}\sum_{t=1}^{T}\max_{i\in S_{t}}\|x_{ti}\|_{H_{t}^{-1}}^{2},$
where the second inequality holds by Lemma D.1. Combining the upper bound for the terms (A) and (B), with probability at least $1-\delta$, we have
@@
$+10\beta_{T}(\delta)^{2}\sum_{t=1}^{T}\max_{i\in S_{t}}\|x_{ti}\|_{H_{t}^{-1}}^{2}.$ (D.3)
Now, we bound each term of (D.3) respectively. For the first term, we decompose it as follows:
@@
$+\sum_{t=1}^{T}\sum_{i\in S_{t}}p_{t}(i|S_{t},\mathbf{w}^{\star})\left(p_{t}(0|S_{t},\mathbf{w}^{\star})-p_{t}(0|S_{t},\mathbf{w}_{t+1})\right)\|x_{ti}\|_{H_{t}^{-1}}.$ (D.4)
To bound the first term on the right-hand side of (D.4), we apply the Cauchy-Schwarz inequality.
@@
$\leq\frac{\sqrt{v_{0}K}}{(v_{0}+Ke^{-1})}\sqrt{T\cdot 2d\log\left(1+\frac{T}{d\lambda}\right)},$ (D.5)
where the last inequality holds by Lemma D.2.
Now, we bound the second term on the right-hand side of (D.4). Let the virtual context for the outside option be $x_{t0}=\mathbf{0}$. Then, by the mean value theorem, there exists $\bm{\xi}_{t}=(1-c)\mathbf{w}^{\star}+c\mathbf{w}_{t+1}$ for some $c\in(0,1)$ such that
@@
$\leq\frac{4d}{\kappa}\beta_{T}(\delta)\log\left(1+\frac{T}{d\lambda}\right)$ (D.6)
where the last inequality holds by Lemma D.2.
Finally, we bound the third term on the right-hand side of (D.4). By the mean value theorem, there exists $\bm{\xi}^{\prime}_{t}=(1-c^{\prime})\mathbf{w}^{\star}+c^{\prime}\mathbf{w}_{t+1}$ for some $c^{\prime}\in(0,1)$ such that
@@
$\leq\frac{2d}{\kappa}\beta_{T}(\delta)\log\left(1+\frac{T}{d\lambda}\right),$ (D.7)
where the last inequality holds by Lemma D.2. By plugging (D.5), (D.6), and (D.7) into (D.4) and multiplying $2\beta_{T}(\delta)$, we get
@@
$\leq 2\sqrt{2}\frac{\sqrt{v_{0}K}}{(v_{0}+Ke^{-1})}\beta_{T}(\delta)\sqrt{dT}\sqrt{\log\left(1+\frac{T}{d\lambda}\right)}+\frac{12d}{\kappa}\beta_{T}(\delta)^{2}\log\left(1+\frac{T}{d\lambda}\right).$ (D.8)
Moreover, by applying Lemma D.2, we can directly bound the second term of (D.3).
$10\beta_{T}(\delta)^{2}\sum_{t=1}^{T}\max_{i\in S_{t}}\|x_{ti}\|_{H_{t}^{-1}}^{2}\leq 10\beta_{T}(\delta)^{2}\cdot\frac{2}{\kappa}d\log\left(1+\frac{T}{d\lambda}\right).$ (D.9)
Finally, plugging (D.8) and (D.9) into (D.3), we obtain
@@
D.2 Useful Lemmas for Theorem 2
D.2.1 Proof of Lemma D.1
Proof of Lemma D.1.

### v6

@@ D.3 Proofs of Lemmas for Theorem 1
$=\frac{v_{0}K\left(\exp\left(x^{\top}\mathbf{w}_{V}\right)-\exp\left(\hat{x}^{\top}\mathbf{w}_{V}\right)\right)}{\left(v_{0}+K\exp\left(x^{\top}\mathbf{w}_{V}\right)\right)\left(v_{0}+K\exp\left(\hat{x}^{\top}\mathbf{w}_{V}\right)\right)}$
$\geq\frac{v_{0}K\left(\exp\left(x^{\top}\mathbf{w}_{V}\right)-\exp\left(\hat{x}^{\top}\mathbf{w}_{V}\right)\right)}{(v_{0}+Ke)^{2}},$ (D.5)
where the inequality holds since $\max\left\{\exp\left(x^{\top}\mathbf{w}_{V}\right),\exp\left(\hat{x}^{\top}\mathbf{w}_{V}\right)\right\}\leq e$. To further bound the right-hand side of  (D.5), we use the fact that $21+a\leq e^{a}\leq 1+a+a^{2}/$ 2 for all $a\in[0,1]$, which can be easily shown by Taylor expansion. Thus, we get
@@
where the last inequality holds because $(\sqrt{d}\epsilon)^{2}\leq\delta\epsilon/\sqrt{d}$ when $\epsilon\in(0,1/d\sqrt{d})$. This concludes the proof. ∎
D.3.2 Proof of Lemma D.2
Proof of Lemma D.2.
Fix a round $t$, an assortment $\tilde{S}_{t}$, and $\tilde{U}_{t}$. Let $U=\tilde{U}_{t}$. Define $m_{j}(\tilde{S}_{t}):=\sum_{x_{U}\in\tilde{S}_{t}}\mathbf{1}\{j\in U\}/K$. Let $\{p_{i}\}_{i\in\tilde{S}_{t}\cup\{0\}}$ and $\{q_{i}\}_{i\in\tilde{S}_{t}\cup\{0\}}$ be the probabilities of choosing item $i$ under parameterizations $\mathbf{w}_{V}$ and $\mathbf{w}_{V\cup\{j\}}$, respectively. Then, we have
$\operatorname{KL}\left(\mathbb{P}_{V}(\cdot|\tilde{S}_{t})\|\mathbb{P}_{V\cup\{j\}}(\cdot|\tilde{S}_{t})\right)=\sum_{i\in\tilde{S}_{t}\cup\{0\}}p_{i}\log\frac{p_{i}}{q_{i}}\leq\sum_{i\in\tilde{S}_{t}\cup\{0\}}p_{i}\frac{p_{i}-q_{i}}{q_{i}}\leq\sum_{i\in\tilde{S}_{t}\cup\{0\}}\frac{(p_{i}-q_{i})^{2}}{q_{i}},$
@@
$\leq C_{0}\cdot\frac{v_{0}K^{2}}{(v_{0}+K)^{3}}\cdot\frac{m_{j}(\tilde{S}_{t})\epsilon^{2}}{d},$ (D.6)
@@
$\sum_{i\in\tilde{S}_{t},j\notin U}\frac{(p_{i}-q_{i})^{2}}{q_{i}}=0,$ (D.7)
@@
Recall that for any $i\in\tilde{S}_{t}$, $q_{i}\geq\frac{e^{-1}}{v_{0}+Ke}$. Then, for any $i\in\tilde{S}_{t}$ corresponding to $x_{i}=\hat{x}$ and $j\in U$, we have
@@
$\leq C_{1}\cdot\frac{v_{0}^{2}K}{(v_{0}+K)^{3}}\cdot\frac{m_{j}(\tilde{S}_{t})\epsilon^{2}}{d},$ (D.8)
@@
Combining (D.6), (D.7), and (D.8), we derive that
@@
$\operatorname{KL}\left(P_{V}\|Q_{V\cup\{j\}}\right)$ $=\sum_{t=1}^{T}\mathbb{E}_{V}\left[\operatorname{KL}\left(\mathbb{P}_{V}(\cdot|\tilde{S}_{t})\|\mathbb{P}_{V\cup\{j\}}(\cdot|\tilde{S}_{t})\right)\right]$
@@
where the equality holds by the chain rule of relative entropy (cf. Exercise 14.11 of Lattimore and Szepesvári [31]). This concludes the proof. ∎
Appendix E Proof of Theorem 2
@@
E.1 Main Proof of Theorem 2
Before presenting the proof, we introduce useful lemmas, whose proof can be found in Appendix E.2. Lemma E.1 shows the optimistic utility for the context vectors.
Lemma E.1.
Let $\alpha_{ti}=x_{ti}^{\top}\mathbf{w}_{t}+\beta_{t}(\delta)\|x_{ti}\|_{H_{t}^{-1}}$. If $\mathbf{w}^{\star}\in\mathcal{C}_{t}(\delta)$, then we have
@@ E.1 Main Proof of Theorem 2
Lemma E.3 is a $K$-free elliptical potential lemma that improves upon the one presented in Lemma 10 of Perivier and Goyal [43] in terms of $K$. Lemma 10 of Perivier and Goyal [43] states: $\sum_{s=1}^{t}\sum_{i\in S_{s}}p_{s}(i|S_{s},\mathbf{w}^{\star})p_{s}(0|S_{s},\mathbf{w}^{\star})\|x_{si}\|_{H_{s}(\mathbf{w}^{\star})^{-1}}^{2}\leq 2dK\log\left(\lambda_{t+1}+\frac{2tK}{d}\right)$ and $\sum_{s=1}^{t}\max_{i\in S_{s}}\|x_{si}\|_{H_{s}(\mathbf{w}^{\star})^{-1}}^{2}\leq 2d\left(K+\frac{1}{\kappa}\right)\log\left(\lambda_{t+1}+\frac{2tK}{d}\right)$, where $H_{t}(\mathbf{w})=\sum_{s=1}^{t-1}\mathcal{G}_{s}(\mathbf{w})+\lambda_{t}\mathbf{I}_{d}$.
Lemma E.2.
@@
Moreover, we provide a tighter bound for the second derivative of the expected revenue than that presented in Lemma 12 of Perivier and Goyal [43]. Lemma 12 of Perivier and Goyal [43] states: $\left|\frac{\partial^{2}Q}{\partial i\partial j}\right|\leq 5$.
Lemma E.3.
Define $Q:\mathbb{R}^{K}\rightarrow\mathbb{R}$, such that for any $\mathbf{u}=(u_{1},\dots,u_{K})\in\mathbb{R}^{K}$, $Q(\mathbf{u})=\sum_{i=1}^{K}\frac{\exp(u_{i})}{v_{0}+\sum_{k=1}^{K}\exp(u_{k})}$. Let $p_{i}(\mathbf{u})=\frac{\exp(u_{i})}{v_{0}+\sum_{k=1}^{K}\exp(u_{k})}$. Then, for all $i\in[K]$, we have
@@
First, we bound the regret as follows:
$\sum_{t=1}^{T}$ $R_{t}(S_{t}^{\star},\mathbf{w}^{\star})-R_{t}(S_{t},\mathbf{w}^{\star})=\sum_{t=1}^{T}\left[\sum_{i\in S_{t}^{\star}}p_{t}(i|S_{t}^{\star},\mathbf{w}^{\star})-\sum_{i\in S_{t}}p_{t}(i|S_{t},\mathbf{w}^{\star})\right]$
@@
$\leq\sum_{t=1}^{T}\left[\frac{\sum_{i\in S_{t}}\exp(\alpha_{ti})}{v_{0}+\sum_{j\in S_{t}}\exp(\alpha_{tj})}-\frac{\sum_{i\in S_{t}}\exp(x_{ti}^{\top}\mathbf{w}^{\star})}{v_{0}+\sum_{j\in S_{t}}\exp(x_{tj}^{\top}\mathbf{w}^{\star})}\right]=\sum_{t=1}^{T}\tilde{R}_{t}(S_{t})-R_{t}(S_{t},\mathbf{w}^{\star}),$
where the first inequality holds by Lemma E.1, and the last inequality holds by the assortment selection of Algorithm 1.
Now, we define $Q:\mathbb{R}^{K}\rightarrow\mathbb{R}$, such that for all $\mathbf{u}=(u_{1},\dots,u_{K})^{\top}\in\mathbb{R}^{K}$, $Q(\mathbf{u})=\sum_{i=1}^{K}\frac{\exp(u_{i})}{v_{0}+\sum_{j=1}^{K}\exp(u_{j})}$. Noting that $S_{t}$ always contains $K$ elements since the expected revenue is an increasing function in the uniform reward setting, we can write $S_{t}=\{i_{1},\dots,i_{K}\}$. Moreover, for all $t\geq 1$, let $\mathbf{u}_{t}=(u_{ti_{1}},\dots u_{ti_{K}})^{\top}=(\alpha_{ti_{1}},\dots,\alpha_{ti_{K}})^{\top}$ and $\mathbf{u}^{\star}_{t}=(u_{ti_{1}}^{\star},\dots u_{ti_{K}}^{\star})^{\top}=(x_{ti_{1}}^{\top}\mathbf{w}^{\star},\dots,x_{ti_{K}}^{\top}\mathbf{w}^{\star})^{\top}$. Then, by a second order Taylor expansion, we have
@@
$=\underbrace{\sum_{t=1}^{T}\nabla Q(\mathbf{u}_{t}^{\star})^{\top}(\mathbf{u}_{t}-\mathbf{u}^{\star}_{t})}_{\texttt{(A)}}+\underbrace{\frac{1}{2}\sum_{t=1}^{T}(\mathbf{u}_{t}-\mathbf{u}^{\star}_{t})^{\top}\nabla^{2}Q(\bar{\mathbf{u}}_{t})(\mathbf{u}_{t}-\mathbf{u}^{\star}_{t})}_{\texttt{(B)}},$ (E.1)
@@
$\leq 2\beta_{T}(\delta)\sum_{t=1}^{T}\sum_{i\in S_{t}}p_{t}(i|S_{t},\mathbf{w}^{\star})p_{t}(0|S_{t},\mathbf{w}^{\star})\|x_{ti}\|_{H_{t}^{-1}},$ (E.2)
where the first inequality holds by Lemma E.1, and the last inequality holds because $\beta_{t}(\delta)$ is increasing for $t\in[T]$.
@@
$\leq\sum_{t=1}^{T}\sum_{i\in S_{t}}\sum_{j\in S_{t},j\neq i}|u_{ti}-u_{ti}^{\star}|p_{i}(\bar{\mathbf{u}}_{t})p_{j}(\bar{\mathbf{u}}_{t})|u_{tj}-u_{tj}^{\star}|+\frac{3}{2}\sum_{t=1}^{T}\sum_{i\in S_{t}}(u_{ti}-u_{ti}^{\star})^{2}p_{i}(\bar{\mathbf{u}}_{t}),$ (E.3)
where the inequality is by Lemma E.3. To bound the first term in (E.3), by applying the AM-GM inequality, we get
@@
$\leq\sum_{t=1}^{T}\sum_{i\in S_{t}}(u_{ti}-u_{ti}^{\star})^{2}p_{i}(\bar{\mathbf{u}}_{t}).$ (E.4)
By plugging (E.4) into (E.3), we have
@@
$\leq 10\beta_{T}(\delta)^{2}\sum_{t=1}^{T}\max_{i\in S_{t}}\|x_{ti}\|_{H_{t}^{-1}}^{2},$ (E.5)
where the second inequality holds by Lemma E.1. Combining the upper bound for the terms (A) and (B), with probability at least $1-\delta$, we have
@@
$+10\beta_{T}(\delta)^{2}\sum_{t=1}^{T}\max_{i\in S_{t}}\|x_{ti}\|_{H_{t}^{-1}}^{2}.$ (E.6)
Now, we bound each term of (E.6) respectively. For the first term, we decompose it as follows:
@@
$+\sum_{t=1}^{T}\sum_{i\in S_{t}}p_{t}(i|S_{t},\mathbf{w}^{\star})\left(p_{t}(0|S_{t},\mathbf{w}^{\star})-p_{t}(0|S_{t},\mathbf{w}_{t+1})\right)\|x_{ti}\|_{H_{t}^{-1}}.$ (E.7)
To bound the first term on the right-hand side of (E.7), we apply the Cauchy-Schwarz inequality.
@@
$\leq\frac{\sqrt{v_{0}K}}{(v_{0}+Ke^{-1})}\sqrt{T\cdot 2d\log\left(1+\frac{T}{d\lambda}\right)},$ (E.8)
where the last inequality holds by Lemma E.2.
Now, we bound the second term on the right-hand side of (E.7). Let the virtual context for the outside option be $x_{t0}=\mathbf{0}$. Then, by the mean value theorem, there exists $\bm{\xi}_{t}=(1-c)\mathbf{w}^{\star}+c\mathbf{w}_{t+1}$ for some $c\in(0,1)$ such that
@@
$\leq\frac{4d}{\kappa}\beta_{T}(\delta)\log\left(1+\frac{T}{d\lambda}\right)$ (E.9)
where the last inequality holds by Lemma E.2.
Finally, we bound the third term on the right-hand side of (E.7). By the mean value theorem, there exists $\bm{\xi}^{\prime}_{t}=(1-c^{\prime})\mathbf{w}^{\star}+c^{\prime}\mathbf{w}_{t+1}$ for some $c^{\prime}\in(0,1)$ such that
@@
$\leq\frac{2d}{\kappa}\beta_{T}(\delta)\log\left(1+\frac{T}{d\lambda}\right),$ (E.10)
where the last inequality holds by Lemma E.2. By plugging (E.8), (E.9), and (E.10) into (E.7) and multiplying $2\beta_{T}(\delta)$, we get
@@
$\leq 2\sqrt{2}\frac{\sqrt{v_{0}K}}{(v_{0}+Ke^{-1})}\beta_{T}(\delta)\sqrt{dT}\sqrt{\log\left(1+\frac{T}{d\lambda}\right)}+\frac{12d}{\kappa}\beta_{T}(\delta)^{2}\log\left(1+\frac{T}{d\lambda}\right).$ (E.11)
Moreover, by applying Lemma E.2, we can directly bound the second term of (E.6).
$10\beta_{T}(\delta)^{2}\sum_{t=1}^{T}\max_{i\in S_{t}}\|x_{ti}\|_{H_{t}^{-1}}^{2}\leq 10\beta_{T}(\delta)^{2}\cdot\frac{2}{\kappa}d\log\left(1+\frac{T}{d\lambda}\right).$ (E.12)
Finally, plugging (E.11) and (E.12) into (E.6), we obtain
@@
Remark E.1.
If the boundedness assumption on the parameter is relaxed to $\|\mathbf{w}\|_{2}\leq B$, since $\beta_{t}(\delta)=\mathcal{O}\left(B\sqrt{d}\log t\log K\right)$, we have $\textup{{Reg}}_{T}(\mathbf{w}^{\star})=\tilde{\mathcal{O}}\left(Be^{B}\frac{\sqrt{v_{0}K}}{v_{0}+K}d\sqrt{T}+\frac{1}{\kappa}d^{2}\right).$ It’s important to note that one of our main goals is to explicitly demonstrate the regret depends on $K$ and $v_{0}$. In deriving such a result, the dependence on $e^{B}$ is unavoidable to our best knowledge. Note that for non-uniform rewards, the regret bound does not depend on $e^{B}$ (refer Remark H.1).
E.2 Proofs of Lemmas for Theorem 2
E.2.1 Proof of Lemma E.1
Proof of Lemma E.1.